# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v4

@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{split}\bm{{\Sigma}_{Z+\epsilon}^{-1}}=&(\bm{U}+\sigma^{2}_{\epsilon}\bm{uu}^{T})^{-1}\\ =&\bm{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sigma^{2}_{\epsilon}\bm{u}^{T}\bm{U}^{-1}\bm{u}}\bm{U}^{-1}\bm{u}\bm{u}^{T}\bm{U}^{-1}\\ =&\bm{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}\bm{U}^{-1}\bm{1}_{k}\bm{U}^{-1}\\ =&\lambda\begin{bmatrix}\frac{1}{\lambda\sigma^{2}_{Z_{1}}}-\frac{1}{\sigma^{4}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{2}}}-\frac{1}{\sigma^{4}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k}}}\\ \vdots&\vdots&&\vdots&\vdots\\ -\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{2}}}&...&\frac{1}{\lambda\sigma^{2}_{Z_{k-1}}}-\frac{1}{\sigma^{4}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{k-1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{k}}}-\frac{1}{\sigma^{4}_{Z_{k}}}\\ \end{bmatrix}\end{split}$ (40)
@@
where $\bm{U}^{-1}=\mathrm{diag}((\sigma^{2}_{Z_{1}})^{-1},...,(\sigma^{2}_{Z_{k}})^{-1})$ and $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}$.
@@
Therefore, substitute Equation 40 into ${|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$, we can obtain:
$\begin{split}&{|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\sigma^{2}_{Y_{k}}\end{bmatrix}-\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\bm{{\Sigma}_{Z+\epsilon}^{-1}}\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\mathrm{cov}^{2}(Y_{1},Z_{1}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{1}}}-\frac{\lambda}{\sigma^{4}_{Z_{1}}})&...&\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ \vdots&&\vdots\\ \mathrm{cov}(Y_{k},Z_{k}+\epsilon)\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&...&\sigma^{2}_{Y_{k}}-\mathrm{cov}^{2}(Y_{k},Z_{k}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{k}}}-\frac{\lambda}{\sigma^{4}_{Z_{k}}})\\ \end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&&\\ &\ddots&\\ &&\sigma^{2}_{Y_{k}}-\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}+\lambda\begin{bmatrix}\frac{1}{\sigma^{4}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{1},Z_{1})\mathrm{cov}(Y_{k},Z_{k})\\ \vdots&&\vdots\\ \frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{k},Z_{k})\mathrm{cov}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{4}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}\right|\end{split}$
@@
We use the notation $\bm{v}=\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}^{T}$, and $\bm{V}=\mathrm{diag}(\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1}),\cdots,\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k}))$. And utilize the rule of determinants of sums Marcus (1990), then we have:
$\begin{split}\left|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}\Sigma^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}\right|=&\ \left|(\bm{\Sigma_{Y}-V})+\lambda\bm{v}\bm{v}^{T}\right|\\ =&\ \left|\bm{\Sigma_{Y}-V}\right|+\lambda\bm{v}^{T}(\bm{\Sigma_{Y}-V})^{-1}\bm{v}.\end{split}$ (42)
@@
where $(\bm{\Sigma_{Y}-V})^{*}$ is the adjoint of the matrix $(\bm{\Sigma_{Y}-V})$. For simplicity, we can rewrite ${|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$ as:
@@
$\begin{split}&\left|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}\Sigma^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}\right|\\ =&\prod_{i=1}^{k}\left(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\cdot\frac{1}{\sigma^{2}_{Z_{i}}}\right)+\Omega\end{split}$ (43)
@@
where $\Omega=\lambda\cdot\bm{v}^{T}\cdot\bm{(\Sigma_{Y}-V)}^{*}\cdot\bm{v}$. The specific value of $\Omega$ can be obtained as:
@@
$\begin{split}\Omega=\lambda\cdot\begin{bmatrix}\dfrac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\dfrac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\cdot\begin{bmatrix}V_{11}&0&0\\ 0&\ddots&0\\ 0&0&V_{kk}\end{bmatrix}\cdot\begin{bmatrix}\dfrac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})\\ \vdots\\ \dfrac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\end{split}$ (44)
@@
$\begin{split}\Omega=&\lambda\sum_{i=1}^{k}\frac{\frac{1}{\sigma^{4}_{Z_{i}}}\mathrm{cov}^{2}(Y_{i},Z_{i})\prod_{j=1}^{k}(\sigma^{2}_{Y_{j}}-\mathrm{cov}^{2}(Y_{j},Z_{j})\frac{1}{\sigma^{2}_{Z_{j}}})}{(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})}\\ =&\lambda\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}\\ \end{split}$ (46)
@@
$\begin{split}&{|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})\end{split}$ (47)
Accordingly, $|\bm{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|$ is:
$|\bm{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|=\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))$ (48)
As a result, $\frac{|\bm{\Sigma_{Y|Z+\epsilon}}|}{|\bm{\Sigma_{Y|Z}}|}$ is expressed as:
$\frac{|\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Y|Z+\epsilon}}|}=\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})}$ (49)
Combine Equations 49 and 39 together, the entropy change is expressed as:
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})}\end{split}$ (50)
It is difficult to tell that Equation 50 is greater or smaller than 00 directly. But one thing for sure is that when there is no Gaussian noise, Equation 50 equals 0. However, we can use another way to compare the numerator and denominator in Equation 50. Instead, we use the symbol $M$ to compare the numerator and denominator using subtraction. Let:
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\bm{\epsilon})>0&\mathrm{if}\ M>0\\ \bigtriangleup S(\mathcal{T},\bm{\epsilon})\leq 0&\mathrm{if}\ M\leq 0\end{matrix}\right.$ (53)
From the above equations, the sign of the entropy change is determined by the statistical properties of the embeddings and labels. Since ${\epsilon}^{2}\geq 0$, $\lambda\geq 0$ and $\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}\geq 0$, we need to have a deep dive into the residual part, i.e.,
@@
where $\rho_{Z_{i}Y_{i}}$ is the correlation coefficient, and $\rho^{2}_{Z_{i}Y_{i}}\in[0,1]$. Eq. 54 is greater than 0, As a result, the sign of the entropy change in the Gaussian noise case is negative. We can conclude that Gaussian noise added to the latent space is harmful to the task.

### v5

@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{split}\boldsymbol{{\Sigma}_{Z+\epsilon}^{-1}}=&(\boldsymbol{U}+\sigma^{2}_{\epsilon}\boldsymbol{uu}^{T})^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sigma^{2}_{\epsilon}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\boldsymbol{u}}\boldsymbol{U}^{-1}\boldsymbol{u}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}\boldsymbol{U}^{-1}\boldsymbol{1}_{k}\boldsymbol{U}^{-1}\\ =&\lambda\begin{bmatrix}\frac{1}{\lambda\sigma^{2}_{Z_{1}}}-\frac{1}{\sigma^{4}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{2}}}-\frac{1}{\sigma^{4}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k}}}\\ \vdots&\vdots&&\vdots&\vdots\\ -\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{2}}}&...&\frac{1}{\lambda\sigma^{2}_{Z_{k-1}}}-\frac{1}{\sigma^{4}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{k-1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{k}}}-\frac{1}{\sigma^{4}_{Z_{k}}}\\ \end{bmatrix}\end{split}$ (39)
@@
where $\boldsymbol{U}^{-1}=\mathrm{diag}((\sigma^{2}_{Z_{1}})^{-1},...,(\sigma^{2}_{Z_{k}})^{-1})$ and $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}$.
@@
Therefore, substitute Equation 39 into $|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|$}, we can obtain:
[…]
@@
We use the notation $\boldsymbol{v}=\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}^{T}$, and $\boldsymbol{V}=\mathrm{diag}(\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1}),\cdots,\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k}))$. And utilize the rule of determinants of sums [36], then we have:
$\begin{split}|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|=&|(\boldsymbol{\Sigma_{Y}-V})+\lambda\boldsymbol{v}\boldsymbol{v}^{T}|\\ =&|\boldsymbol{\Sigma_{Y}-V}|+\lambda\boldsymbol{v}^{T}\boldsymbol{(\boldsymbol{\Sigma_{Y}-V})^{*}}\boldsymbol{v}\end{split}$ (41)
@@
where $(\boldsymbol{\Sigma_{Y}-V})^{*}$ is the adjoint of the matrix $(\boldsymbol{\Sigma_{Y}-V})$. For simplicity, we can rewrite $|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|$} as:
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})+\Omega\end{split}$ (42)
@@
where $\Omega=\lambda\boldsymbol{v}^{T}\boldsymbol{(\boldsymbol{\Sigma_{Y}-V})}^{*}\boldsymbol{v}$. The specific value of $\Omega$ can be obtained as:
@@
$\begin{split}\Omega=\lambda\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\begin{bmatrix}V_{11}&&\\ &\ddots&\\ &&V_{kk}\end{bmatrix}\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})\\ \vdots\\ \frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\end{split}$ (43)
@@
[…]
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})\end{split}$ (46)
Accordingly, $|\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|$ is:
$|\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|=\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))$ (47)
As a result, $\frac{|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}{|\boldsymbol{\Sigma_{Y|Z}}|}$ is expressed as:
$\frac{|\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}=\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})}$ (48)
Combine Equations 48 and 38 together, the entropy change is expressed as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})}\end{split}$ (49)
It is difficult to tell whether Equation 49 is greater or smaller than 0 directly. But one thing is for sure that when there is no Gaussian noise, Equation 49 equals 0. However, we can use another way to compare the numerator and denominator in Equation 49. Instead, we use the symbol $M$ to compare the numerator and denominator using subtraction. Let:
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&\mathrm{if}\ M>0\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&\mathrm{if}\ M\leq 0\end{matrix}\right$. (52)
From the above equations, the sign of the entropy change is determined by the statistical properties of the embeddings and labels. Since ${\epsilon}^{2}\geq 0$ and $\lambda\geq 0$, we need to have a deep dive into the residual part, i.e.,
@@
where $\rho_{Z_{i}Y_{i}}$ is the correlation coefficient, and $\rho^{2}_{Z_{i}Y_{i}}\in[0,1]$. Eq. 53 is greater than 0, As a result, the sign of the entropy change in the Gaussian noise case is less than zero. We can conclude that Gaussian noise added to the latent space is harmful to the task.